Grade 9Mechanics

A body freely falling from rest has a velocity V after it falls through a height h.The distance it has to fall further for its velocity to become double is ?

Parth
7 Years agoGrade 9

1 Answer

Arun
7 Years ago

w^2 = v ^2 + 2g s

s is the distance to find.

w = 2V

Therefore

s = 3v^2/(2g )

v^2 = 2gH

Therefore

s = 3*2g H /(2g) = 3 H